# Scalable Multi-Task Learning for Particle Collision Event Reconstruction with Heterogeneous Graph Neural Networks

### v1

@@ Abstract
The growing luminosity frontier at the Large Hadron Collider is challenging the reconstruction and analysis of particle collision events. Increased particle multiplicities are straining latency and storage requirements at the data acquisition stage, while new complications are emerging, including higher background levels and more frequent particle vertex misassociations. This in turn necessitates the development of more holistic and scalable reconstruction methods that take advantage of recent advances in machine learning. We propose a novel Heterogeneous Graph Neural Network (HGNN) architecture featuring unique representations for diverse particle collision relationships and integrated graph pruning layers for scalability. Trained with a multi-task paradigm in an environment mimicking the LHCb experiment, this HGNN significantly improves beauty hadron reconstruction performance. Notably, it concurrently performs particle vertex association and graph pruning within a single framework. We quantify reconstruction and pruning performance, demonstrate enhanced inference time scaling with event complexity, and mitigate potential performance loss using a weighted message passing scheme.
@@ 1 Introduction
Over the past two decades, the field of neural networks has witnessed rapid advancements, driving breakthroughs across diverse domains such as natural language processing, image analysis, and scientific computation. Architectures ranging from convolutional neural networks [1] to the more recent transformer [2] models have played pivotal roles in these developments. Meanwhile, graph neural networks [3] (GNNs) have emerged as a powerful tool for representing complex datasets of variable-size that lack explicit spatial or sequential structures, effectively modeling interactions among multiple entities and their interconnections. Building on this success, heterogeneous GNNs [4] (HGNNs) extend conventional GNNs by incorporating multiple types of nodes and edges, enabling richer and more context-specific representations for complex, multi-relational data. HGNNs find extensive application in areas such as recommendation systems (modeling users, items, and interactions) [5], bioinformatics (gene-disease-drug networks) [6], and natural language processing tasks like cross-lingual text classification [7].
@@
Concurrently, in particle physics recent trends have shown a rising employment of machine learning (ML) across several domains including simulation, detector reconstruction and particle classification [9, 10]. HGNNs and multi-task learning are particularly applicable at particle collider experiments given that particle collision events are inherently heterogeneous involving several object types and their relationships with one another. However, despite their potential, HGNNs have rarely been applied in particle physics, where most applications of GNNs rely on homogeneous GNN architectures [11, 12, 13]. Unifying HGNNs and multi-task learning allows for several reconstruction tasks in parallel.
@@
The LHCb experiment [14] at CERN’s Large Hadron Collider focuses on high-precision studies of beauty ($b$) and charm ($c$) quarks. Its main goals are to test the Standard Model (SM) through measurements of CP violation, rare beauty hadron decays, and flavor-changing neutral current processes that are highly sensitive to new physics beyond the SM. During Runs 1 and 2 (2010-2018), LHCb recorded $9\,\mathrm{fb}^{-1}$ of proton-proton collision data at a luminosity of $4\times 10^{32}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$. A recent upgrade (Upgrade I) for Runs 3 and 4 increased the luminosity to $2\times 10^{33}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$, resulting in an average of 5 collisions per event and a charged particle track multiplicity of around 150 [15]. Looking ahead, Upgrade II at the High-Luminosity LHC is expected to boost the luminosity tenfold, yielding roughly 50 collisions per event [16] and a track multiplicity of around 1,000 particles.
@@
The rising collision event complexity at the LHCb experiment presents significant challenges for data acquisition and measurement precision. While LHCb’s Upgrade I trigger framework [17] already transitioned to partial event storage, moving beyond the traditional approach of storing full events which was feasible when disk space was less constrained [18], the challenges intensify under HL-LHC conditions. The anticipated increase in particle multiplicity and the common occurrence of multiple heavy-hadron decays within single events will further strain storage resources. Consequently, refining strategies to systematically identify and preserve the most valuable event components remains an important objective, especially considering the diverse heavy-hadron species and decays per species, whose combinatorics raise to tens of thousands of different possible decays of interest. While exclusive selections are advantageous for storage and crucial for many physics analyses to avoid biases, capturing sufficient information for studies requiring context from the underlying event (e.g., for flavour tagging or semileptonic decay spectroscopy) presents an ongoing challenge. Developing techniques that can selectively preserve this broader event information, when needed for specific physics goals, while adhering to stringent latency constraints ($\mathcal{O}$(100ms) per event) and storage limitations ($\mathcal{O}$(10PB) per year), will be beneficial for maximising the physics potential.
@@
Another critical issue emerging in the context of increased luminosity conditions at LHCb is primary vertex (PV) misassociation. PVs correspond to the proton-proton interaction points, and under high-luminosity conditions, multiple such interactions can occur in a single event. Misassociation arises when tracks or decay products from overlapping proton-proton collisions are incorrectly attributed to the wrong PV. This can severely degrade vertex resolution and bias the measurement of key observables, such as the $B$-hadron decay flight distance and direction, ultimately affecting the precision of determinations of time-dependent CP violation [19] and measurements of decays with missing energy [20, 21]. Addressing PV misassociation is therefore paramount to maintaining the high-precision performance required for LHCb’s physics program in the high-luminosity era, prompting the need for innovative reconstruction algorithms and advanced machine learning techniques.
@@
Recently, we proposed the Deep Full Event Interpretation (DFEI) algorithm [22], employing GNNs to perform a multi-stage inclusive reconstruction of beauty hadrons in LHCb collision events. Despite its potential, the computational cost and scalability of its multi-stage approach pose significant challenges with regards to meeting the low-latency requirements of a real-time trigger environment. Furthermore, the algorithm does not address the pressing issue of PV misassociation. Building upon the success of our earlier work, we propose a HGNN architecture with integrated graph pruning for scalability, which is trained with multiple objectives to perform beauty hadron reconstruction and PV association within a unified framework.
@@ 2 Related work
Particle decay reconstruction typically follows an exclusive strategy, in which final‑state particles are successively combined into intermediate structures to form complete decay chains for specific topologies. This process proceeds sequentially, combining final-state particles into higher-level structures based on predefined decay modes and kinematic constraints, such as invariant mass thresholds or conservation of momentum and energy. At each level, particle identification techniques and machine learning classification algorithms may be applied to improve the reconstruction accuracy and resolve ambiguities.
@@
Beyond traditional methods that target individual decay channels, intermediate strategies exist that bridge these and the fully holistic reconstruction we propose. Key examples include the LHCb topological trigger [23, 24] and the Belle II tag-side reconstruction algorithm known as the Full Event Interpretation (FEI) [25]. The LHCb topological trigger identifies beauty hadron decays based on predefined characteristic topologies, which rely on multivariate classifiers trained on an ensemble of decay modes making its selection inherently guided by predefined exclusive states. Similarly, the FEI performs a hierarchical reconstruction of a large number of beauty decay chains with a dedicated multivariate classifier for each unique particle decay.
@@
More recently, there have been several efforts towards a fully inclusive reconstruction of beauty hadron decays with GNNs at LHCb and Belle II [26, 27, 22]. These developments were prompted by Kahn et al. [26] with the introduction of a novel edge classification target for hierarchical decay chains known as the Lowest Common Ancestor Generations (LCAG) matrix. This compact representation enables one to learn the hierarchical structure of a decay solely from its final-state particles. For each edge relation between final-state particles a multi-class label is used, which signifies the generational class of the shared ancestor.
@@
Our previous publication [22], the Deep Full Event Interpretation (DFEI), expanded on this work demonstrating the inclusive reconstruction of beauty hadron decays with GNNs within the hadronic environment of LHCb, which is complicated by a large number of background particles. To overcome these difficulties, we employed a multi-stage approach. First, a node-pruning GNN filters out background nodes based on kinematic and topological features. Next, an edge-pruning GNN removes unlikely associations, simplifying the event graph. Lastly, a final GNN processes the remaining graph and performs the edge classification of LCAG values between nodes, enabling a separation and hierarchical reconstruction of multiple possible beauty hadron decay chains in each event.
@@
Beyond particle decay reconstruction, GNNs have seen significant adoption in particle physics [11]. For charged particle tracking, GNNs are employed to connect detector hits (nodes) into particle trajectories by classifying potential track segments (edges) [28]. GNNs have demonstrated an improved performance for flavor tagging of beauty hadron events at Belle II with GFlaT [29]. GNNs are also increasingly used for jet classification and reconstruction tasks, representing jets as point clouds or graphs of constituent particles to distinguish between different originating particles and better reconstruct the kinematical quantities of the jet [30, 31, 32]. Furthermore, GNNs are employed for particle flow (PF) algorithms, aiming to provide an end-to-end ML approach that combines information from different subdetectors to reconstruct a complete list of particles [33]. While most applications use homogeneous GNNs, HGNNs have been applied to improve hadronic $\tau$ lepton identification by treating tracks and energy clusters as distinct node types within a jet graph [12] and in novel designs for track reconstruction that explicitly account for different detector sensor types (e.g., pixel vs. strip hits) [13].
@@
To manage computational costs and focus on relevant relations, two main strategies are employed: graph pruning and dynamic graph construction. Graph pruning typically starts with a larger, often geometrically constrained graph, and then removes edges deemed unlikely to represent true physical connections. This is common in tracking pipelines (e.g. ExaTrkX), where initial filtering steps or GNN-based edge classifiers prune the graph significantly before track finding [34, 35]. Usually thresholding the output scores of edge-classifying GNNs serves as the pruning mechanism. Alternatively, dynamic graph construction methods adapt the graph connectivity during the learning process. Techniques like k-Nearest Neighbors (k-NN) applied in a learned latent space allow the graph structure to evolve, connecting nodes based on their learned features rather than just initial geometry. This is exemplified by architectures like EdgeConv (used in ParticleNet for jet tagging) [30] and GravNet (used for calorimeter clustering) [36], which dynamically define edges. Both pruning and dynamic construction aim to improve the overall performance and scalability of GNNs.
@@
Lastly, multi-task learning (MTL), the paradigm we adopt, has become increasingly relevant in particle physics, where complex analysis often involves inferring multiple correlated properties or performing hierarchical reconstruction steps. For instance, in jet physics, “foundation models” are being developed using MTL principles, pre-training on large datasets for self-supervised tasks like jet generation or masked particle prediction, then fine-tuning for various downstream tasks like jet tagging or property prediction, aiming to create a universal jet representation [37, 38, 39]. In neutrino physics event reconstruction, MTL is used in GNNs like NuGraph2 to simultaneously perform tasks such as detector hit classification and semantic segmentation (assigning hits to particle types like electron or muon) [40]. The concept also extends to using auxiliary tasks to aid a primary goal, such as predicting track origins alongside flavour tagging in ATLAS’s GNN jet tagger (GN1) [31]. Beyond particle physics, MTL is also used in related fields such as fusion energy for real-time plasma equilibrium reconstruction, simultaneously predicting multiple plasma parameters and profile distributions [41].
@@ 3.1 Graph Neural Networks
GNNs encompass a diverse range of architectures, each tailored to specific tasks and graph structures. Key approaches include Graph Convolutional Networks (GCNs) [42], which extend traditional convolutions by aggregating and transforming features from connected nodes, and Graph Attention Networks (GATs) [43], which use attention mechanisms to weight node contributions. Message Passing Neural Networks (MPNNs) [44] further unify these methods under a general framework where interconnected nodes iteratively exchange and update messages to learn graph representations.
In this paper we build upon the GNN introduced by Battaglia et al. [45], which provides a more comprehensive and versatile framework for MPNNs learning representations at multiple levels, including nodes, edges, and globally, making it well-suited to capture the hierarchical and relational information inherent in particle collisions. Equation 1 summarises the update equations for the GNN, which consist of edge, node and graph updates.
@@
The edge update function $\phi^{e}$ is a learnable Multi-Layer-Perceptron (MLP), which takes as input the existing edge representation $e^{k}$ for edge $k$, adjoining receiver and sender node representations ($v^{r_{k}}$ and $v^{s_{k}}$) and a global representation $u$. The subsequent node update with MLP $\phi^{v}$ for node $i$, takes as input the node $v^{i}$ representation, the edge to node aggregation $\bar{e}^{i}$ and $u$. With $\bar{e}^{i}$ aggregating edge representations for the set of edges $E^{\prime i}$ that node $i$ receives using a aggregation function $\rho^{e\rightarrow v}$. Lastly, sets of edges and nodes, $E^{\prime}$ and $V^{\prime}$, are aggregated globally yielding, $\bar{e}^{\prime}$ and $\bar{v}^{\prime}$, using aggregation functions $\rho^{e\rightarrow u}$ and $\rho^{v\rightarrow u}$. The global update $u^{\prime}$ with MLP $\phi^{u}$ takes as input the global aggregations and the existing global representation, $u$.

### v2

@@ Abstract
The growing luminosity frontier at the Large Hadron Collider is challenging the reconstruction and analysis of particle collision events. Increased particle multiplicities are straining latency and storage requirements at the data acquisition stage, while new complications are emerging, including higher background levels and more frequent particle vertex misassociations. This in turn necessitates the development of more holistic and scalable reconstruction methods that take advantage of recent advances in machine learning. We propose a novel Heterogeneous Graph Neural Network (HGNN) architecture featuring unique representations for diverse particle collision relationships and integrated graph pruning layers for scalability. Trained with a multi-task paradigm in an environment mimicking the LHCb experiment, this HGNN significantly improves the beauty hadron reconstruction performance. Notably, it concurrently performs particle vertex association and graph pruning within a single framework. We quantify the reconstruction and pruning performance, demonstrate enhanced inference time scaling with event complexity, and mitigate potential performance loss using a weighted message passing scheme.
@@ 1 Introduction
Over the past two decades, the field of neural networks has witnessed rapid advancements, driving breakthroughs across diverse domains such as natural language processing, image analysis, and scientific computation. Architectures ranging from convolutional neural networks [1] to more recent transformer [2] models have played pivotal roles in these developments. Meanwhile, graph neural networks [3] (GNNs) have emerged as a powerful tool for representing complex datasets of variable size that lack explicit spatial or sequential structures, effectively modeling interactions among multiple entities and their interconnections. Building on this success, heterogeneous GNNs [4] (HGNNs) extend conventional GNNs by incorporating multiple types of nodes and edges, enabling richer and more context-specific representations for complex, multi-relational data. HGNNs find extensive applications in areas such as recommendation systems (modeling users, items, and interactions) [5], bioinformatics (gene-disease-drug networks) [6], and natural language processing tasks such as cross-lingual text classification [7].
@@
Concurrently, in particle physics recent trends have shown an increase in the application of machine learning (ML) across several domains including simulation, detector reconstruction and particle classification [9, 10]. HGNNs and multi-task learning are particularly suited to particle collider experiments given that particle collision events are inherently heterogeneous involving several object types and their relationships with one another. However, despite their potential, HGNNs have rarely been applied in particle physics [11, 12], where most applications rely on homogeneous GNN architectures [13]. Unifying HGNNs and multi-task learning allows several reconstruction tasks to be performed in parallel.
@@
The LHCb experiment [14] at CERN’s Large Hadron Collider focused on high-precision studies of beauty ($b$) and charm ($c$) quarks. Its main goal is to test the Standard Model (SM) through measurements of CP violation, rare beauty-hadron decays, and flavor-changing neutral current processes that are highly sensitive to new physics beyond the SM. During Runs 1 and 2 (2010-2018), LHCb recorded $9\,\mathrm{fb}^{-1}$ of proton-proton collision data at a luminosity of $4\times 10^{32}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$. A recent upgrade (Upgrade I) for Runs 3 and 4 increased the luminosity to $2\times 10^{33}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$, resulting in an average of five collisions per event and a charged particle track multiplicity of approximately 150 [15]. Looking ahead, the recently approved LHCb Upgrade II at the High-Luminosity LHC is expected to boost the luminosity tenfold, yielding on average 50 collisions per event [16] and a track multiplicity of around 1,000 particles.
@@
The increasing collision event complexity at the LHCb experiment presents significant challenges for data acquisition and measurement precision. While LHCb’s Upgrade I trigger framework [17] has already transitioned to partial event storage, moving beyond the traditional approach of storing full events that was feasible when the disk space was less constrained [18], the challenges will intensify under Upgrade II conditions. The anticipated increase in particle multiplicity and the common occurrence of multiple heavy-hadron decays within single events will further strain storage resources. Consequently, refining strategies to systematically identify and preserve the most valuable event components remains an important objective, especially when considering the diverse heavy-hadron species and decays per species, whose combinatorics amount to tens of thousands of different possible decays of interest. While exclusive selections are advantageous for storage, capturing sufficient information for studies requiring context from the underlying event (e.g., for flavour tagging or measurements of semileptonic decays) presents an ongoing challenge. Developing techniques that can selectively preserve this broader event information, when needed for specific physics goals, while adhering to stringent latency constraints ($\mathcal{O}$(100ms) per event and CPU core) and storage limitations ($\mathcal{O}$(10PB) per year), will be beneficial for maximising the physics potential of LHCb.
@@
Another critical issue emerging in the context of increased luminosity conditions at the LHCb is primary vertex (PV) misassociation. PVs correspond to proton-proton interaction points, and under high-luminosity conditions, multiple such interactions can occur in a single event. PV misassociation arises when tracks or decay products from overlapping proton-proton collisions are incorrectly attributed to a PV. This can severely degrade the PV resolution and bias the measurement of key observables, such as the beauty-hadron decay flight distance and direction, ultimately affecting the precision of determinations of time-dependent CP violation [19] and measurements of decays with missing energy [20, 21]. Therefore, addressing PV misassociation is paramount to maintaining the high-precision performance required for LHCb’s physics program in the high-luminosity era, prompting the need for innovative reconstruction algorithms and advanced machine learning techniques.
@@
Recently, we proposed the Deep Full Event Interpretation (DFEI) algorithm [22], which employs GNNs to perform a multi-stage inclusive reconstruction of beauty hadrons in LHCb collision events. Despite its potential, the computational cost and scalability of its multi-stage approach pose significant challenges with regard to meeting the low-latency requirements of a real-time trigger environment. Furthermore, the algorithm did not address the pressing issue of PV misassociation. Building on the success of our earlier work, we propose an HGNN architecture with integrated graph pruning for scalability, which is trained with multiple objectives to perform beauty hadron reconstruction and PV association within a unified framework.
@@ 2 Related work
Particle decay reconstruction typically follows an exclusive strategy in which final‑state particles are successively combined into intermediate structures to form complete decay chains for specific topologies. This process proceeds sequentially by combining final-state particles into higher-level structures based on predefined decay modes and kinematic constraints, such as invariant mass thresholds or the conservation of momentum and energy. At each level, particle identification techniques and machine learning classification algorithms can be applied to improve reconstruction accuracy and resolve ambiguities.
@@
In addition to traditional methods that target individual decay channels, intermediate strategies exist that bridge these and the fully holistic reconstruction we propose. Key examples include the LHCb topological trigger [23, 24] and the Belle II tag-side reconstruction algorithm known as Full Event Interpretation (FEI) [25]. The LHCb topological trigger identifies beauty-hadron decays based on predefined characteristic topologies, which rely on multivariate classifiers trained on an ensemble of decay modes, making their selection inherently guided by predefined exclusive states. Similarly, FEI performs a hierarchical reconstruction of a large number of beauty-hadron decay chains with a dedicated multivariate classifier for each unique particle decay.
@@
More recently, several efforts have been made towards a fully inclusive reconstruction of beauty-hadron decays with GNNs at LHCb and Belle II [26, 27, 22]. These developments were prompted by Kahn et al. [26] with the introduction of a novel edge classification target for hierarchical decay chains known as the lowest common ancestor generations (LCAG) matrix. This compact representation enables one to learn the hierarchical structure of a decay solely from its final-state particles. For each edge relation between final-state particles, a multi-class label is used, which signifies the generational class of the shared ancestor.
@@
In our previous publication [22], the Deep Full Event Interpretation (DFEI), we expanded on this work demonstrating the inclusive reconstruction of beauty-hadron decays with GNNs within the hadronic environment of LHCb, which is complicated by a large number of background particles. To overcome these difficulties, we employed a multi-stage approach. First, a node-pruning GNN filters out background nodes based on kinematic and topological features. Next, an edge-pruning GNN removes unlikely associations and simplifies the event graph. Finally, a GNN processes the remaining graph and performs the edge classification of LCAG values, enabling the separation and hierarchical reconstruction of multiple possible beauty-hadron decay chains in each event.
@@
Beyond particle decay reconstruction, GNNs have seen significant adoption in particle physics [13]. For charged-particle tracking, GNNs are employed to connect detector hits (nodes) into particle trajectories by classifying potential track segments (edges) [28]. GNNs have demonstrated improved performance for flavor tagging of beauty-hadron events at Belle II with GFlaT [29]. GNNs are also increasingly used for jet classification and reconstruction tasks, representing jets as point clouds or graphs of constituent particles to distinguish between different originating particles and to better reconstruct the kinematic quantities of the jet [30, 31, 32]. Furthermore, GNNs are employed for particle flow (PF) algorithms, which aim to provide an end-to-end ML approach that combines information from different subdetectors to reconstruct a complete list of particles [33]. While most applications use homogeneous GNNs, HGNNs have been applied to improve hadronic $\tau$ lepton identification by treating tracks and energy clusters as distinct node types within a jet graph [11] and in novel designs for track reconstruction that explicitly account for different detector sensor types (e.g., pixel vs. strip hits) [12].
@@
To manage computational costs and focus on relevant relations, two main strategies are employed: graph pruning and dynamic graph construction. Graph pruning typically starts with a larger, often geometrically constrained graph, and then removes edges deemed unlikely to represent true physical connections. This is common in tracking pipelines (e.g. ExaTrkX), where initial filtering steps or GNN-based edge classifiers prune the graph significantly before track finding [34, 35]. Usually thresholding the output scores of edge-classifying GNNs serves as the pruning mechanism. Alternatively, dynamic graph construction methods can adapt the graph connectivity during the learning process. Techniques such as k-nearest neighbors (k-NN) applied in a learned latent space allow the graph structure to evolve, connecting nodes based on their learned representations. This is exemplified by architectures such as EdgeConv (used in ParticleNet for jet tagging) [30] and GravNet (used for calorimeter clustering) [36], which dynamically define edges. Both pruning and dynamic construction aim to improve the overall performance and scalability of GNNs.
@@
Finally, multi-task learning (MTL), the paradigm we adopt, has become increasingly relevant in particle physics, where complex analysis often involves inferring multiple correlated properties or performing hierarchical reconstruction steps. For instance, in jet physics, “foundation models” are being developed using MTL principles, pre-training on large datasets for self-supervised tasks such as jet generation or masked particle prediction, and then fine-tuning for various downstream tasks such as jet tagging or property prediction, with the aim of creating a universal jet representation [37, 38, 39]. In neutrino physics event reconstruction, NuGraph2 applied MTL with GNNs to simultaneously perform tasks such as detector hit classification and semantic segmentation (assigning hits to electron or muon particle types) [40]. The concept also extends to using auxiliary tasks to aid a primary goal, such as predicting track origins alongside jet flavour in ATLAS’s GNN jet tagger (GN1) [31]. Beyond particle physics, MTL is also used in related fields such as fusion energy for real-time plasma equilibrium reconstruction, simultaneously predicting multiple plasma parameters and profile distributions [41].
@@ 3.1 Graph Neural Networks
GNNs encompass a diverse range of architectures, each tailored to specific tasks and graph structures. Key approaches include Graph Convolutional Networks (GCNs) [42], which extend traditional convolutions to graphs by aggregating and transforming features from connected nodes, and Graph Attention Networks (GATs) [43], which use attention mechanisms to weight node contributions. Message Passing Neural Networks (MPNNs) [44] further unify these methods under a general framework where interconnected nodes iteratively exchange and update messages to learn graph representations.
In this paper we build upon the GNN introduced by Battaglia et al. [45], which provides a more comprehensive and versatile framework for MPNNs learning representations at multiple levels, including nodes, edges, and globally, making it well suited to capture the hierarchical and relational information inherent in particle collisions. Equation 1 summarises the update equations for the GNN, which consist of edge, node and graph updates.
@@
The edge update function $\phi^{e}$ is a learnable multi-layer-perceptron (MLP), that takes as input the existing edge representation $e^{k}$ for edge $k$, adjoining receiver and sender node representations ($v^{r_{k}}$ and $v^{s_{k}}$) and a global representation $u$. The subsequent node update with MLP $\phi^{v}$ for node $i$, takes as input the node $v^{i}$ representation, the edge-to-node aggregation $\bar{e}^{\prime i}$ and $u$. Here, $\bar{e}^{\prime i}$ aggregates edge representations for the set of edges $E^{\prime i}$ that node $i$ receives using an aggregation function $\rho^{e\rightarrow v}$. Finally, sets of edges and nodes, $E^{\prime}$ and $V^{\prime}$, are aggregated globally to give $\bar{e}^{\prime}$ and $\bar{v}^{\prime}$ using the aggregation functions $\rho^{e\rightarrow u}$ and $\rho^{v\rightarrow u}$. The global update $u^{\prime}$ with MLP $\phi^{u}$ takes as input global aggregations of edges ($\bar{e}^{\prime}$) and nodes ($\bar{v}^{\prime}$) and the existing global representation, $u$.